# Bit of a Close Talker: A Practical Guide to Serverless Cloud Co-Location Attacks

### v1

@@ V-A Attack Construction Guide
    Focus on one specific choice of the corresponding scheduling parameter, and accurately target the victim’s function. This occurs when this choice of scheduling parameter can be known by an attacker. For example, in a package-aware scheduler [8], when a victim hosts a popular serverless service like data analysis, the used packages are usually predictable. This tends to be the most efficient way of launching attacks, as shown by our results in Section VI.
@@
Step 4. In Step 4, we observe whether $\mathrm{F}_{2}$ (Auto-Scaling) is involved. Auto-scaling tends to happen when a serverless function is in high demand. The scheduler will assign more function hosts in the system, causing the function host to spread in the cloud system. An attacker can exploit this feature to increase the physical server coverage of a single function by creating repeated invocation bursts, which is similar to the attack method used in [47].
@@ V-B Constructed Attacks for Selected Scheduling Algorithms
    Targeting the auto-scaling feature (Helper), in this attack strategy attackers stress one of their own serverless functions by creating a burst of invocations, aiming to exploit the auto-scaling mechanism to achieve co-location.
@@ 3. M3-1
    Targeting PASch, the attacker exploits the locality-based scheduling algorithm by creating functions with the same package dependencies as the victim function, increasing the likelihood of co-location with victim function hosts.
@@ 4. M3-2
We categorize M1, M2, and M3-2 as scatter-based attacks, since these attacks achieve co-location by scattering function hosts to increase the coverage, while M3-1 only targets the correct physical server. It is worth noting that M3-2 and M1 are different, although both involve creating multiple functions and deploying them. In M3-2, the attacker has identified the exploitable locality feature and can hence use this feature to scatter instances, while in M1 the attacker simply creates multiple copies of a single function and all parameters remain the same.
@@ VI-A Serverless System Implementation
Besides conducting experiments in the aforementioned simulator, we also deploy serverless cloud systems with different schedulers to a cluster. Due to resource and cost constraints, we choose to evaluate on a $50$-node cluster on CloudLab [16] consisting of $50$ r320 instances. Each node has one 8-core E5-2450 CPU and 16GB Memory. We use one node as the central scheduler, and the others as workers.
The implementation of the serverless platform follows Abdi et al. [1]. We implement the scheduling and function invocation functionalities entirely in the Dask Distributed framework [14], which is a Python-based distributed computing framework. Our implementation involves approximately 500 lines of changes across two files that together comprise around 16,000 lines of Python code. The changes are distributed as follows: modifications to scheduler.py for the scheduling algorithm implementation, and updates to worker.py to adapt the system for serverless operations. Functions are executed through the Azure Functions Host runtime [35]. Upon cold-start, the server that is scheduled to execute the function will pull the function code from a remote destination and invoke the Azure Functions Host runtime. We use function benchmarks from benchmark sets such as FunctionBench [25]. It is worth noting that this research focuses completely on the scheduler side, hence we can safely omit environment configuration and execution details during the function execution process. We will use the log files produced by the scheduler node and worker nodes to obtain function placement information and calculate the required metrics. The four selected algorithms are implemented in this system.
@@ VI-B Metrics
Simulation Experiments. The simulator enables us to flexibly change experiment parameters, e.g. cluster size and conduct a large number of experiments. We launch 50 short-duration experiments in a 1000-node simulated cluster and use the percentage of experiments where the attacker successfully co-locates with the victim as our metric.
Cloud Experiments. Unlike in the traditional cloud instance placement situation where instances are relatively fixated [18, 17], the flexible and scalable nature of serverless scheduling complicates the co-location measurements. Simply counting how many of the invocations to victim functions are co-located with attacker functions does not take the constant termination and start of instances into account. For example, an attacker’s function invocation causes the corresponding function host to be placed together with a victim, resulting in subsequent invocations to this victim function being risky. The success of these subsequent invocations should be attributed to the specific function invocation that leads to co-location and shouldn’t be counted multiple times.
@@
  • Attacker Efficiency (AE), i.e. the percentage of attacker function instance placement that results in co-location with victim instances;
  • Placement Accuracy (PA), i.e. the percentage of victim instances that co-locate with attack instances.
@@ VI-C Results
We perform the evaluation of our attack in both simulation and a cluster. In the simulation experiments, the metric of focus is the short-term success rate, since it is easier to launch massive experiments in simulation. In each experiment, for a relatively short duration, we specify a victim user, and the attacker launches attack functions. We then collect the generated log files and calculate the percentage of successful experiments. In cluster experiments, we focus on a more realistic attack process with a longer duration and obtain the AE and PA data to measure the attack efficiency.
@@
From Fig. 6, we can see that in both PASch and OpenWhisk, the attackers are able to achieve relatively high co-location performance as they increase the number of attack functions. Also, it can be seen from the results of M3-2 that even if attackers do not possess full knowledge of victim package dependencies, they can still achieve co-location by scattering their function instances. Below are some additional findings.
@@
By observing the two curves in Fig. 6 regarding PASch (the scheduler that considers package locality), we can see that the attacker is able to reach almost $100\%$ success rate. This indicates that as long as the locality feature of the scheduler is identified by the attacker, the attacker is able to easily achieve a high co-location success rate. If the user-submitted information is known by the attacker, e.g. the user utilizes certain packages, an attacker can precisely target the victim user’s function host in the cluster by providing identical package requirements, which is the M3-1 attack strategy we provided in Section V. Otherwise, the attacker can construct multiple functions with various settings to ensure coverage, which is the M3-2 strategy. Either way, the attacker is capable of achieving a high success rate.
@@
Cluster Experiments. We also conduct experiments in a 50-node CloudLab [16] cluster. All functions are invoked continuously during the experiments, with function invocation being submitted every two seconds. It takes 3-6 hours to run one experiment. We select to report AE and PA results of these experiments, which reflect the efficiency of attacker’s attempts and the effects of attacks on victim function hosts. The results on the CloudLab cluster [16] are shown in Fig. 7. We can observe that the PASch scheduler and the Random scheduler have the highest PA values, indicating that most user’s placed instances are co-located with a victim instance during its lifetime. By looking at the two curves corresponding to the Helper Scheduler (Helper: M1 vs Helper: M2), we can conclude that compared to using one attack function and invoking repeatedly to trigger the auto-scaling policies of the scheduler, using multiple functions results in higher PA values, i.e. more user function hosts are co-located with attack function hosts.
@@
It is worth noting that as we only specify a single victim and only repeat our experiments a few times, unlike the simulation experiments that are repeated hundreds of times, we can see random fluctuations in the results, which is close to the nature of this task in the real world. However, we are still able to see that with a relatively small number of attack functions, the attacker can achieve co-location with $\sim 50\%$ victim function hosts in most schedulers.
@@
Finding 3: Attack efficiency of scatter-based attacks (i.e. M1, M2, M3-2) are relatively low.
From Fig. 7 (a), we can see that most of the AE values of scatter-based attacks are below $0.1$ (OpenWhisk: M1, Helper: M1, and PASch: M3-2), meaning only a small portion of those placed attack function hosts achieve co-location with victim function hosts. The only exception is the Random scheduler, where the M1 attack strategy achieves an AE value of 1. An explanation of this is provided in Appendix B. Overall, our observations indicate that attacks would be more efficient if locality-optimization features of schedulers could be identified and the attacker could obtain related information about the victim.
The results in both simulation and a cluster show that our constructed attack can enable attackers to effectively achieve co-location with victim function hosts. We also show that the most important thing an attacker can do is to identify scheduling locality features, as exploiting these features can greatly increase the success rate, as well as improve attack efficiency.
Transferrability. We further examine whether the attack methods are transferrable. We apply attack methods in Section V to mismatched types of schedulers. The experiments are conducted in the aforementioned simulator. The results are presented in Fig. 8.
@@
The PASch [8]-related curves in Fig. 8 show that it is hard to achieve co-location if attackers do not exploit the locality features. The success rates of both curves (PASch: M1 and PASch: M2) remain constant at 0. This is due to the fact that this type of scheduler tends to restrict the placement of functions to specific locations, and without targeting the correct scheduling features the attacker will not be able to identify the victim’s location.
Results from our simulation show that generally, without targeting correct scheduling features, an attacker will not be able to achieve co-location even with massive function invocations. The only exception is the scheduler with auto-scaling features, where it is more efficient to deploy multiple functions instead of simply triggering the auto-scaling policy to scale out. The results show that it is therefore important to identify exploitable features of schedulers, i.e. scheduler fingerprinting.
@@ VII-A Scheduler Fingerprinting
Compared to Phase 1, we can see that in Phase 2 the Azure Functions scheduler selects a completely different set of physical servers to execute function hosts, indicating that the function start locations are irrelevant to previous placement results. Besides, in Phase 2, the two constructed functions each cover 4 physical machines, showing that the scheduler of Azure Functions does not have the tendency to place different functions from the same user on the same physical machines.
@@ VII-B Co-Location Attack
We create two accounts, one as the victim and the other as the attacker. The victim user runs one function in Azure Functions, which is our target. The attacker creates 64 different function applications, each containing one function. Both users utilize the Azure Consumption Plan, and the functions created by both users are the fingerprinting function so that we can verify co-location. We deploy these functions, invoke them, and analyze the collected traces to determine if co-location is achieved.
@@ VIII-A Double-Dip Scheduler
We propose Double-Dip, a scheduling algorithm that aims to provide “soft” server isolation for different users. The goal of Double-Dip is to minimize overlap of physical servers between different users. Double-Dip works as follows:
@@
The pseudocode of the Double-Dip scheduling algorithm is provided as the following:
@@
Integrating Double-Dip into commercial serverless clouds requires only minimal modifications to existing scheduling pipelines, as it operates entirely at the scheduling layer and leaves execution mechanisms unchanged. Major platforms such as AWS Lambda, Azure Functions, and Google Cloud Run already maintain per-function metadata; Double-Dip simply leverages this to keep a lightweight active-user map per host and apply a user-aware rule during placement, and remains fully compatible with modern elastic environments. The scheduling overhead is small, consisting of constant-time user checks and a single pass to identify the least-user-variety host—operations already common in production load balancers. Our cost experiments further show that this logic scales well: even as we increase users and requests by an order of magnitude, ’s warm-start behavior remains close to Helper and above OpenWhisk, with overhead saturating once workers have seen all function types.
@@ VIII-B Evaluation
We conducted our experiments in a simulated environment to evaluate the performance of Double-Dip scheduling algorithm in a serverless cloud setting, comparing it with Helper. The simulation setup consisted of 100 nodes, each with the capacity to handle up to 1024 functions. The environment featured one victim function and a variable number of attackers, ranging from 5 to 50. Each attacker was programmed to perform 10 attack attempts during the simulation. To ensure statistical significance and reduce variability, each result was obtained by averaging data from 1000 independent simulation runs.
@@ VIII-C Cost Evaluation
Figure 12 reports the warm-start ratios for all configurations. Across all settings, we observe that Helper consistently has the highest warm-start ratio, followed by Double-Dip. This behavior is expected: Helper aggressively reuses recently active nodes, maximizing warm-starts but providing no user-level isolation. Double-Dip lies in between, yet it still retains most of the warm-start advantages of Helper while offering substantially stronger isolation guarantees. The lower warm-start ratios in the smallest configuration reflect the system’s rapid scale-up, where new workers are activated quickly; this behavior is expected in serverless platforms and further highlights that Double-Dip scales smoothly as the cluster expands.
@@ IX-A Security Implications in Scheduler Design
Cloud scheduler is an important part of cloud systems. They are crucial to the performance of the whole system and at the same time the first barrier to cloud attackers. A good scheduler should optimize performance by fully exploiting function locality features while simultaneously ensuring unpredictability in scheduling results to enhance security.
Our results help provide some insights into the security aspect of serverless scheduler design. On the one hand, we show that pure random schedulers are susceptible to co-location attacks, and attackers are guaranteed to achieve a high success rate as long as they prepare a sufficient number of attack functions to deploy. On the other hand, optimization policies that result in more deterministic instance placements, e.g. package-aware scheduling, can also be exploited by attackers to achieve co-location with victims. These results inspire a hybrid approach, where the scheduler still preserves the original optimization technology which leads to determinism in instance placement, but at the same time adds randomness to the system to avoid the scheduling results being easily identified. The added randomness should not be excessive as well, otherwise the attacker would still be able to scatter attack function hosts and achieve co-location with victim hosts. It is important for the algorithm designer to strike the balance point and ensure a performant and secure system.
Besides, it is also important to provide user isolation. The co-location vulnerability in this paper largely originates from the fact that different users share the same scheduling pool, i.e. the whole cluster. If different users’ scheduling pools are isolated or only partially overlapped, the difficulty of achieving co-location would significantly increase. This should not influence the performance of users’ functions, since each function only scales out to a small number of physical servers even under a large burst of invocation requests, as shown by the results from Azure Functions.
IX-B Other Potential Defence
Although the deployment of Double-Dip algorithm proposed in this paper can significantly reduce the co-location threat of serverless clouds, it does not provide 100% guarantee that the attacker and victim will not be co-located. Given the nature of serverless system design, i.e. it tries to abstract away cloud management details from users and increase server utilization rate from the vendor side, the only solution to guarantee instance isolation is to turn to traditional Infrastructure-as-a-Service (IaaS) clouds and reserve dedicated servers to use. However, this will significantly increase the cost. Below we outline a ballpark estimation to compare the cost of using dedicated servers to provide isolation guarantee and Double-Dip.
Since the changes we propose in the Double-Dip algorithm is oblivious to users, we assume the customers are still charged the same price as before. We calculate using the following setting:

### v3

@@ V-A Attack Construction Guide
    Focus on one specific choice of the corresponding scheduling parameter, and accurately target the victim’s function. This occurs when the choice of scheduling parameter can be known by an attacker. For example, in a package-aware scheduler [8], when a victim hosts a popular serverless service like data analysis, the used packages are usually predictable. This tends to be the most efficient way of launching attacks, as shown by our results in Section VI.
@@
Step 4. In Step 4, we observe whether $\mathrm{F}_{2}$ (Auto-Scaling) is involved. Auto-scaling tends to happen when a serverless function is in high demand. The scheduler will assign more function hosts in the system, causing the function hosts to spread across the cloud system. An attacker can exploit this feature to increase the physical server coverage of a single function by creating repeated invocation bursts, which is similar to the attack method used in [47].
@@ V-B Constructed Attacks for Selected Scheduling Algorithms
    Targeting the auto-scaling feature (Helper), in this attack strategy, attackers stress one of their own serverless functions by creating a burst of invocations, aiming to exploit the auto-scaling mechanism to achieve co-location.
@@ 3. M3-1
    Targeting PASch, the attacker exploits the locality-based scheduling algorithm by creating functions with the same package dependencies as the victim function, increasing the likelihood of co-location with the victim function hosts.
@@ 4. M3-2
We categorize M1, M2, and M3-2 as scatter-based attacks, since these attacks achieve co-location by scattering function hosts to increase the coverage, while M3-1 only targets the correct physical server. It is worth noting that M3-2 and M1 are different, although both involve creating multiple functions and deploying them. In M3-2, the attacker has identified the exploitable locality feature and can hence use this feature to scatter instances, while in M1, the attacker simply creates multiple copies of a single function, and all parameters remain the same.
@@ VI-A Serverless System Implementation
Besides conducting experiments in the aforementioned simulator, we also deploy serverless cloud systems with different schedulers to a cluster. Due to resource and cost constraints, we choose to evaluate on a $50$-node cluster on CloudLab [16] consisting of $50$ r320 instances. Each node has one 8-core E5-2450 CPU and 16GB of Memory. We use one node as the central scheduler, and the others as workers.
The implementation of the serverless platform follows Abdi et al. [1]. We implement the scheduling and function invocation functionalities entirely in the Dask Distributed framework [14], which is a Python-based distributed computing framework. Our implementation involves approximately 500 lines of changes across two files that together comprise around 16,000 lines of Python code. The changes are distributed as follows: modifications to scheduler.py for the scheduling algorithm implementation, and updates to worker.py to adapt the system for serverless operations. Functions are executed through the Azure Functions Host runtime [35]. Upon cold-start, the server that is scheduled to execute the function will pull the function code from a remote destination and invoke the Azure Functions Host runtime. We use function benchmarks from benchmark sets such as FunctionBench [25]. It is worth noting that this research focuses completely on the scheduler side; hence, we can safely omit environment configuration and execution details during the function execution process. We will use the log files produced by the scheduler node and worker nodes to obtain function placement information and calculate the required metrics. The four selected algorithms are implemented in this system.
@@ VI-B Metrics
Simulation Experiments. The simulator enables us to flexibly change experiment parameters, e.g., cluster size, and conduct a large number of experiments. We launch 50 short-duration experiments in a 1000-node simulated cluster and use the percentage of experiments where the attacker successfully co-locates with the victim as our metric.
Cloud Experiments. Unlike in the traditional cloud instance placement situation, where instances are relatively fixed [18, 17], the flexible and scalable nature of serverless scheduling complicates the co-location measurements. Simply counting how many of the invocations to victim functions are co-located with attacker functions does not take the constant termination and start of instances into account. For example, an attacker’s function invocation causes the corresponding function host to be placed together with a victim, resulting in subsequent invocations to this victim function being risky. The success of these subsequent invocations should be attributed to the specific function invocation that leads to co-location and shouldn’t be counted multiple times.
@@
  • Attacker Efficiency (AE), i.e., the percentage of attacker function instance placement that results in co-location with victim instances;
  • Placement Accuracy (PA), i.e., the percentage of victim instances that co-locate with attack instances.
@@ VI-C Results
We perform the evaluation of our attack in both a simulation and a cluster. In the simulation experiments, the metric of focus is the short-term success rate, since it is easier to launch massive experiments in simulation. In each experiment, for a relatively short duration, we specify a victim user, and the attacker launches attack functions. We then collect the generated log files and calculate the percentage of successful experiments. In cluster experiments, we focus on a more realistic attack process with a longer duration and obtain the AE and PA data to measure the attack efficiency.
@@
From Fig. 6, we can see that in both PASch and OpenWhisk, the attackers are able to achieve relatively high co-location performance as they increase the number of attack functions. Also, it can be seen from the results of M3-2 that even if attackers do not possess full knowledge of the victim package dependencies, they can still achieve co-location by scattering their function instances. Below are some additional findings.
@@
By observing the two curves in Fig. 6 regarding PASch (the scheduler that considers package locality), we can see that the attacker is able to reach almost $100\%$ success rate. This indicates that as long as the locality feature of the scheduler is identified by the attacker, the attacker is able to easily achieve a high co-location success rate. If the user-submitted information is known by the attacker, e.g., the user utilizes certain packages, an attacker can precisely target the victim user’s function host in the cluster by providing identical package requirements, which is the M3-1 attack strategy we provided in Section V. Otherwise, the attacker can construct multiple functions with various settings to ensure coverage, which is the M3-2 strategy. Either way, the attacker is capable of achieving a high success rate.
@@
Cluster Experiments. We also conduct experiments in a 50-node CloudLab [16] cluster. All functions are invoked continuously during the experiments, with function invocation being submitted every two seconds. It takes 3-6 hours to run one experiment. We select to report AE and PA results of these experiments, which reflect the efficiency of the attacker’s attempts and the effects of attacks on the victim function hosts. The results on the CloudLab cluster [16] are shown in Fig. 7. We can observe that the PASch scheduler and the Random scheduler have the highest PA values, indicating that most users’ placed instances are co-located with a victim instance during its lifetime. By looking at the two curves corresponding to the Helper Scheduler (Helper: M1 vs Helper: M2), we can conclude that compared to using one attack function and repeatedly invoking to trigger the auto-scaling policies of the scheduler, using multiple functions results in higher PA values, i.e., more user function hosts are co-located with attack function hosts.
@@
It is worth noting that, as we only specify a single victim and only repeat our experiments a few times, unlike the simulation experiments that are repeated hundreds of times, we can see random fluctuations in the results, which is close to the nature of this task in the real world. However, we are still able to see that with a relatively small number of attack functions, the attacker can achieve co-location with $\sim 50\%$ victim function hosts in most schedulers.
@@
Finding 3: Attack efficiency of scatter-based attacks (i.e. M1, M2, M3-2) is relatively low.
From Fig. 7 (a), we can see that most of the AE values of scatter-based attacks are below $0.1$ (OpenWhisk: M1, Helper: M1, and PASch: M3-2), meaning only a small portion of those placed attack function hosts achieves co-location with victim function hosts. The only exception is the Random scheduler, where the M1 attack strategy achieves an AE value of 1. An explanation of this is provided in Appendix B. Overall, our observations indicate that attacks would be more efficient if locality-optimization features of schedulers could be identified and the attacker could obtain related information about the victim.
The results in both the simulation and a cluster show that our constructed attack can enable attackers to effectively achieve co-location with victim function hosts. We also show that the most important thing an attacker can do is to identify scheduling locality features, as exploiting these features can greatly increase the success rate, as well as improve attack efficiency.
Transferability. We further examine whether the attack methods are transferable. We apply attack methods in Section V to mismatched types of schedulers. The experiments are conducted in the aforementioned simulator. The results are presented in Fig. 8.
@@
The PASch [8]-related curves in Fig. 8 show that it is hard to achieve co-location if attackers do not exploit the locality features. The success rates of both curves (PASch: M1 and PASch: M2) remain constant at 0. This is due to the fact that this type of scheduler tends to restrict the placement of functions to specific locations, and without targeting the correct scheduling features, the attacker will not be able to identify the victim’s location.
Results from our simulation show that generally, without targeting correct scheduling features, an attacker will not be able to achieve co-location even with massive function invocations. The only exception is the scheduler with auto-scaling features, where it is more efficient to deploy multiple functions instead of simply triggering the auto-scaling policy to scale out. The results show that it is therefore important to identify exploitable features of schedulers, i.e., scheduler fingerprinting.
@@ VII-A Scheduler Fingerprinting
Compared to Phase 1, we can see that in Phase 2, the Azure Functions scheduler selects a completely different set of physical servers to execute function hosts, indicating that the function start locations are irrelevant to previous placement results. Besides, in Phase 2, the two constructed functions each cover 4 physical machines, showing that the scheduler of Azure Functions does not have the tendency to place different functions from the same user on the same physical machines.
@@ VII-B Co-Location Attack
We create two accounts, one as the victim and the other as the attacker. The victim user runs one function in Azure Functions, which is our target. The attacker creates 64 different function applications, each containing one function. Both users utilize the Azure Consumption Plan, and the functions created by both users are the fingerprinting function, so that we can verify co-location. We deploy these functions, invoke them, and analyze the collected traces to determine if co-location is achieved.
@@ VIII-A Double-Dip Scheduler
We propose Double-Dip, a scheduling algorithm that aims to provide “soft” server isolation for different users. The goal of Double-Dip is to minimize the overlap of physical servers between different users. Double-Dip works as follows:
@@
The pseudocode of the Double-Dip scheduling algorithm is provided as in Fig. 1.
@@
Integrating Double-Dip into commercial serverless clouds requires only minimal modifications to existing scheduling pipelines, as it operates entirely at the scheduling layer and leaves execution mechanisms unchanged. Major platforms such as AWS Lambda, Azure Functions, and Google Cloud Run already maintain per-function metadata; Double-Dip simply leverages this to keep a lightweight active-user map per host and apply a user-aware rule during placement, and remains fully compatible with modern elastic environments. The scheduling overhead is small, consisting of constant-time user checks and a single pass to identify the least-user-variety host—operations already common in production load balancers. Our cost experiments further show that this logic scales well: even as we increase users and requests by an order of magnitude, warm-start behavior remains close to Helper and above OpenWhisk, with overhead saturating once workers have seen all function types.
@@ VIII-B Evaluation
We conducted our experiments in a simulated environment to evaluate the performance of the Double-Dip scheduling algorithm in a serverless cloud setting, comparing it with Helper. The simulation setup consisted of 100 nodes, each with the capacity to handle up to 1024 functions. The environment featured one victim function and a variable number of attackers, ranging from 5 to 50. Each attacker was programmed to perform 10 attack attempts during the simulation. To ensure statistical significance and reduce variability, each result was obtained by averaging data from 1000 independent simulation runs.
@@ VIII-C Cost Evaluation
Fig. 12 reports the warm-start ratios for all configurations. Across all settings, we observe that Helper consistently has the highest warm-start ratio, followed by Double-Dip. This behavior is expected: Helper aggressively reuses recently active nodes, maximizing warm-starts but providing no user-level isolation. Double-Dip lies in between, yet it still retains most of the warm-start advantages of Helper while offering substantially stronger isolation guarantees. The lower warm-start ratios in the smallest configuration reflect the system’s rapid scale-up, where new workers are activated quickly; this behavior is expected in serverless platforms and further highlights that Double-Dip scales smoothly as the cluster expands.
@@ IX-A Security Implications in Scheduler Design
Cloud scheduler is an important part of cloud systems. They are crucial to the performance of the whole system, and at the same time, the first barrier to cloud attackers. A good scheduler should optimize performance by fully exploiting function locality features while simultaneously ensuring unpredictability in scheduling results to enhance security.
Our results help provide some insights into the security aspect of serverless scheduler design. On the one hand, we show that pure random schedulers are susceptible to co-location attacks, and attackers are guaranteed to achieve a high success rate as long as they prepare a sufficient number of attack functions to deploy. On the other hand, optimization policies that result in more deterministic instance placements, e.g., package-aware scheduling, can also be exploited by attackers to achieve co-location with victims. These results inspire a hybrid approach, where the scheduler still preserves the original optimization technology, which leads to determinism in instance placement, but at the same time adds randomness to the system to avoid the scheduling results being easily identified. The added randomness should not be excessive as well, otherwise the attacker would still be able to scatter attack function hosts and achieve co-location with victim hosts. It is important for the algorithm designer to strike the balance point and ensure a performant and secure system.
Besides, it is also important to provide user isolation. The co-location vulnerability in this paper largely originates from the fact that different users share the same scheduling pool, i.e., the whole cluster. If different users’ scheduling pools are isolated or only partially overlapped, the difficulty of achieving co-location would significantly increase. This should not influence the performance of users’ functions, since each function only scales out to a small number of physical servers even under a large burst of invocation requests, as shown by the results from Azure Functions.
IX-B Other Potential defense
Although the deployment of the Double-Dip algorithm proposed in this paper can significantly reduce the co-location threat of serverless clouds, it does not provide a 100% guarantee that the attacker and victim will not be co-located. Given the nature of serverless system design, i.e., it tries to abstract away cloud management details from users and increase server utilization rate from the vendor side, the only solution to guarantee instance isolation is to turn to traditional Infrastructure-as-a-Service (IaaS) clouds and reserve dedicated servers to use. However, this will significantly increase the cost. Below, we outline a ballpark estimation to compare the cost of using dedicated servers to provide an isolation guarantee and Double-Dip.
Since the changes we propose in the Double-Dip algorithm are oblivious to users, we assume the customers are still charged the same price as before. We calculate using the following settings: